# Prediction of Cellular Malignancy Using Electrical Impedance Signatures and Supervised Machine Learning

### v1

@@ 2.2 Machine Learning Models in Bioelectrical Cancer Classification
Supervised ML algorithms can be trained using datasets compiled from literature and clinical studies to capture diverse patient profiles. In this study, the input features are restricted to key electrical properties -dielectric constant, frequency, conductivity, and characteristic relaxation time constant- as the primary aim is to examine how these properties evolve during the transformation of normal cells to tumorous ones. Understanding these bioimpedance signatures may contribute to future strategies for electrically guided cancer detection or intervention [Hossain2021],[Hossain2024],[Hossain2021b],[Hossain2022],[Hossain2021c]. To improve the model’s robustness, future work will explore the use of simulation tool to generate synthetic data and expand the dataset.
@@ 2.3 Understanding Random Forest Algorithm
Random Forest (RF) is a supervised ML algorithm that constructs an ensemble of decision trees (DT) to improve prediction accuracy and reduce overfitting [Salman2024],[Halabaku2024]. Each decision tree is trained on a random subset of data and selects a random subset of features to make its prediction. In classification tasks, the final output is determined by majority voting among the individual trees, enhancing robustness and generalization of the model. This approach not only enhances prediction accuracy but reduces the risk of overfitting, which is critical when working with biological data that may vary across patients or experimental setups. In the context of this study, RF offers a powerful framework for classifying cells based on their dielectric signatures. Supporting the overarching goal of early, non-invasive cancer detection using bioimpedance characteristics.
@@ 2.4 Understanding Support Vector Machine (SVM)
Another widely used classification model in the biomedical signal analysis is the Support Vector Machine (SVM) [Pisner2020],[Meyer2001]. SVM are particularly effective in high dimensional spaces and are known for their robustness in handling small-medium sized datasets with clear margins of separation. By constructing a hyperplane that best divides the data into classes, SVM can classify dielectric properties of tissues. When data is not linearly separable, often in the case with biological signals, SVM employs kernel functions (e.g. radial basis function, polynomial kernel) to project the data into a higher dimensional feature space where linear separation becomes feasible.
@@ 2.5 Understanding K-Nearest Neighbor
K-Nearest Neighbor (KNN) is a non-parametric instance-based supervised machine learning algorithm that classifies data points based on the classes of their nearest neighbors in the feature space [Bezdek1986],[Zhang2016]. The algorithm relies on a user-defined parameter, k, which determines the number of neighboring points considered when assigning a class label. KNN uses various distance metrics, such as Euclidean distance, Manhattan distance, and Minkowski distance-to compute the similarity between instances [Iqbal2020]. These distance functions are critical in determining the neighborhood structure and can significantly impact classification performance. In the case of breast cancer detection using bioimpedance signatures, KNN can serve as a simple effective baseline model for distinguishing between normal, benign and malignant cells. The algorithm’s strength lies in its simplicity and its adaptability to both linear and nonlinear data distributions without requiring explicit model training. Moreover, it can capture local patterns and is particularly useful in cases where the decision boundary is irregular or data driven. For this study, different values of k will be tested to identify the optimal combination that yields the highest classification accuracy, minimizing false positives or negatives, and improve model reliability. Although KNN can be computationally expensive during inference, especially on large datasets, it remains a useful method for comparative analysis in non-invasive bioimpedance -based cancer classification tasks.
@@ 3 Data Collection and Methodology
For this preliminary study, a sample dataset was compiled from 33 scholarly publications using the keyword “dielectric properties of cancer cells” in academic search engines such as Google Scholar and PubMed Central (PMC). Each manuscript was screened for quantitative data on dielectric properties of normal (healthy), benign and malignant cells. The primary parameters extracted included relative permittivity ($\varepsilon_{r}$), characteristic relaxation time constant ($\tau_{p}$), and conductivity ($\varepsilon$). The frequency range of the datasets spanned from 150 kHz to 20 GHz. Initial analysis revealed that both relative permittivity and conductivity were significantly higher in breast cancer cells compared to their healthy counterparts [Hossain2020],[Hossain2021]. This difference highlights the potential of dielectric properties as discrimination biomarkers for cancer classification. These compiled dielectric parameters were then pre-processed for ML model development. Data from different studies were standardized to ensure consistent units and measurement scales. The resulting standardized datasets was subsequently divided into training and testing subsets, forming the foundation for evaluating the classification performance of various supervised learning algorithms, including RF, SVM and KNN. Three supervised ML algorithms were evaluated using the compiled datasets: RF, SVM and KNN. The performance of these algorithms was assessed using four key metrics: accuracy (ACC), F1 score, recall and precision [Asif2021],[Hossain2025],[Hossain2025b],[Hossain2024_preprint]. Accuracy measures the proportion of correct prediction relative to the total prediction made by the model and is generally more reliable for balanced datasets. However, given the limitations and potential imbalance in the literature -derived datasets, recall and precision provide additional insight. Recall is defined as the ratio of correctly classified actual positives to all actual positives, while precision is the ratio of correctly classified actual positives to all predicted positives [Flach2015]. The F1-score, calculated as the harmonic means of precision and recall, offers a balanced evaluation of both metrics [Diallo2024]. As outlined earlier, the primary objective of this study is to compare the predictive performance of the three models, to determine the pathological status of the breast cell, based on the permittivity and conductivity measurement. The algorithm will be used to optimize performance and the corresponding evaluation metrics will be used to identify the model with the highest predictive capability, thereby reducing the risk of false positives and false negatives. For the RF model, the scikit-learn library was used with default parameters [Pedregosa2011]. The dataset was partitioned into 80% training and 20% testing subsets. The number of estimators varied from 1 to 500, while the maximum tree depth varied from 1 to 10. The corresponding F1 score, and accuracy were recorded, and the results are discussed in the following section. A similar procedure was applied to the SVM and the linear SVC models, also implemented using the scikit-learn library with default parameter (e.g. regularization parameter C=1). The maximum number of iterations varied between 10, 50, 100, 200, 500, 1000 and 2000. The corresponding evaluation metrics were analyzed to investigate the relationship between hyperparameter variation and performance improvement, particularly with respect to the F1 score.
@@ 4 Results and Discussion
Table-I presents the evaluation metrics for the supervised ML models. The RF achieved the highest accuracy of 90% and F1 score =88.3% with a maximum depth of 3 and 4 and 100 estimators. In contrast, the support vector machine (SVM) attains a lower accuracy of 66% and F1 score of 64.4%. Future work will focus on extending the model by incorporating additional parameters, such as capacitance as well as datasets derived from simulation models.

### v2

@@ 2.2 Machine Learning Models in Bioelectrical Cancer Classification
Supervised ML algorithms can be trained using datasets compiled from literature and clinical studies to capture diverse patient profiles. In this study, the input features are restricted to key electrical properties -dielectric constant, frequency, conductivity, and characteristic relaxation time constant- as the primary aim is to examine how these properties evolve during the transformation of normal cells to tumorous ones. Understanding these bioimpedance signatures may contribute to future strategies for electrically guided cancer detection or intervention [9],[57],[58],[59],[60]. To improve the model’s robustness, future work will explore the use of simulation tool to generate synthetic data and expand the dataset.
@@ 2.3 Understanding Random Forest Algorithm
Random Forest (RF) is a supervised ML algorithm that constructs an ensemble of decision trees (DT) to improve prediction accuracy and reduce overfitting [61],[62]. Each decision tree is trained on a random subset of data and selects a random subset of features to make its prediction. In classification tasks, the final output is determined by majority voting among the individual trees, enhancing robustness and generalization of the model. This approach not only enhances prediction accuracy but reduces the risk of overfitting, which is critical when working with biological data that may vary across patients or experimental setups. In the context of this study, RF offers a powerful framework for classifying cells based on their dielectric signatures. Supporting the overarching goal of early, non-invasive cancer detection using bioimpedance characteristics.
@@ 2.4 Understanding Support Vector Machine (SVM)
Another widely used classification model in the biomedical signal analysis is the Support Vector Machine (SVM) [63],[64]. SVM are particularly effective in high dimensional spaces and are known for their robustness in handling small-medium sized datasets with clear margins of separation. By constructing a hyperplane that best divides the data into classes, SVM can classify dielectric properties of tissues. When data is not linearly separable, often in the case with biological signals, SVM employs kernel functions (e.g. radial basis function, polynomial kernel) to project the data into a higher dimensional feature space where linear separation becomes feasible.
@@ 2.5 Understanding K-Nearest Neighbor
K-Nearest Neighbor (KNN) is a non-parametric instance-based supervised machine learning algorithm that classifies data points based on the classes of their nearest neighbors in the feature space [65],[66]. The algorithm relies on a user-defined parameter, k, which determines the number of neighboring points considered when assigning a class label. KNN uses various distance metrics, such as Euclidean distance, Manhattan distance, and Minkowski distance-to compute the similarity between instances [67]. These distance functions are critical in determining the neighborhood structure and can significantly impact classification performance. In the case of breast cancer detection using bioimpedance signatures, KNN can serve as a simple effective baseline model for distinguishing between normal, benign and malignant cells. The algorithm’s strength lies in its simplicity and its adaptability to both linear and nonlinear data distributions without requiring explicit model training. Moreover, it can capture local patterns and is particularly useful in cases where the decision boundary is irregular or data driven. For this study, different values of k will be tested to identify the optimal combination that yields the highest classification accuracy, minimizing false positives or negatives, and improve model reliability. Although KNN can be computationally expensive during inference, especially on large datasets, it remains a useful method for comparative analysis in non-invasive bioimpedance -based cancer classification tasks.
@@ 3 Data Collection and Methodology
For this preliminary study, a sample dataset was compiled from 33 scholarly publications using the keyword “dielectric properties of cancer cells” in academic search engines such as Google Scholar and PubMed Central (PMC). Each manuscript was screened for quantitative data on dielectric properties of normal (healthy), benign and malignant cells. The primary parameters extracted included relative permittivity ($\varepsilon_{r}$), characteristic relaxation time constant ($\tau_{p}$), and conductivity ($\varepsilon$). The frequency range of the datasets spanned from 150 kHz to 20 GHz. Initial analysis revealed that both relative permittivity and conductivity were significantly higher in breast cancer cells compared to their healthy counterparts [2],[9]. This difference highlights the potential of dielectric properties as discrimination biomarkers for cancer classification. These compiled dielectric parameters were then pre-processed for ML model development. Data from different studies were standardized to ensure consistent units and measurement scales. The resulting standardized datasets was subsequently divided into training and testing subsets, forming the foundation for evaluating the classification performance of various supervised learning algorithms, including RF, SVM and KNN. Three supervised ML algorithms were evaluated using the compiled datasets: RF, SVM and KNN. The performance of these algorithms was assessed using four key metrics: accuracy (ACC), F1 score, recall and precision [68],[69],[70],[71]. Accuracy measures the proportion of correct prediction relative to the total prediction made by the model and is generally more reliable for balanced datasets. However, given the limitations and potential imbalance in the literature -derived datasets, recall and precision provide additional insight. Recall is defined as the ratio of correctly classified actual positives to all actual positives, while precision is the ratio of correctly classified actual positives to all predicted positives [72]. The F1-score, calculated as the harmonic means of precision and recall, offers a balanced evaluation of both metrics [73]. As outlined earlier, the primary objective of this study is to compare the predictive performance of the three models, to determine the pathological status of the breast cell, based on the permittivity and conductivity measurement. The algorithm will be used to optimize performance and the corresponding evaluation metrics will be used to identify the model with the highest predictive capability, thereby reducing the risk of false positives and false negatives. For the RF model, the scikit-learn library was used with default parameters [74]. The dataset was partitioned into 80% training and 20% testing subsets. The number of estimators varied from 1 to 500, while the maximum tree depth varied from 1 to 10. The corresponding F1 score, and accuracy were recorded, and the results are discussed in the following section. A similar procedure was applied to the SVM and the linear SVC models, also implemented using the scikit-learn library with default parameter (e.g. regularization parameter C=1). The maximum number of iterations varied between 10, 50, 100, 200, 500, 1000 and 2000. The corresponding evaluation metrics were analyzed to investigate the relationship between hyperparameter variation and performance improvement, particularly with respect to the F1 score.
@@ 4 Results and Discussion
Table2 presents the evaluation metrics for the supervised ML models. The RF achieved the highest accuracy of 90% and F1 score =88.3% with a maximum depth of 3 and 4 and 100 estimators. In contrast, the support vector machine (SVM) attains a lower accuracy of 66% and F1 score of 64.4%. Future work will focus on extending the model by incorporating additional parameters, such as capacitance as well as datasets derived from simulation models.